R130000 is deposited into an account now that pays an interest rate at 10.5%p.a so that an amount of money can be withdrawen from the account every 6 months in perpetuity starting one year from now. If it is decided to increase the value of the withdrawals from the second withdrawal onwards at a rate of 5%p.a then the value of the first withdrawal made one year from now is equal to?
1
Expert's answer
2020-05-25T20:51:56-0400


This is a perpetual annuity:

"130 000\times0.0525+136825\times0.025=6825+3420.63=10245.63"